hartnn (hartnn):

Inverse Laplace transform of \[\log(1+a^2/s^2 )\] There is one property i am not recollecting. I can differentiate the above function in 's' domain and divide the function in 't' time domain by t. Answer : 2 (1-cos(at))/t

hartnn (hartnn):

is this correct ? \(\Large L^{-1}[\dfrac{d}{ds}F(s)] = (-1)^n\dfrac{f(t)}{t}\) ??

hartnn (hartnn):

i am getting that answer with the formula \(\Large L^{-1}[\dfrac{d}{ds}F(s)] = -tf(t)\) wanted to confirm whether its correct or not...

hartnn (hartnn):

Also, can someone verify whether this general formula is correct ot not... \(\Large L^{-1}[\dfrac{d^n}{ds^n}F(s)] = (-1)^n t^nf(t)\)

OpenStudy (neer2890):

Laplace transform of \[t ^{n}f(t)\] \[L[t ^{n}f(t)]=(-1)^{n}. \frac{ d ^{n} }{ ds ^{n} }[F(s)]\]

hartnn (hartnn):

yes, i used that to make that inverse laplace formula i posted...just wanted to be sure its absolutely correct...

OpenStudy (neer2890):

so \[L ^{-1}[(-1)^{n}\frac{ d ^{n} }{ ds ^{n}}[F(s)]]=t ^{n}f(t)\]

hartnn (hartnn):

any references for such inverse lappy formulas ??

OpenStudy (neer2890):

N.P. Bali (Engineering mathematics)

hartnn (hartnn):

is it available online ? ebook ?

OpenStudy (neer2890):

no..,you have to purchase it, it's available for purchasing online

OpenStudy (neer2890):

or the most easy way ,search for laplace transform on google

hartnn (hartnn):

i did... but this is inverse laplace.....wiki didn't gave inverse laplace formulas...

hartnn (hartnn):

*give

OpenStudy (neer2890):

just you have to ajust all laplace transform formulas as i did to make them inverse laplace transform formulas .

hartnn (hartnn):

yes, i did the same. Thanks neer :)

OpenStudy (neer2890):

you are welcome